# Diffusion through complex confining environments: fluctuating triply periodic minimal surfaces

### v1

@@ Abstract
The transport of individual entities through interconnected structures is a process of practical relevance both in biology and technology. Examples are given by diffusive dynamics of molecules in porous structures. In soft environments, this transport can be strongly influenced by fluctuations of the porous structure itself. Here, we focus on triply periodic membrane structures found both in cell organelles and in synthetic amphiphilic systems. We theoretically study the effect of a complex three-dimensional fluctuating environment on the diffusive motion of a test object, using a phase field approach. The rigid spherical test object is energetically forced to not penetrate the membrane. Generally, the pores of the membrane structure can be smaller than the diffusing object. Yet, fluctuations of the membrane can intermittently widen its pores, still allowing for the motion of the larger particles through them. Thus, the object stays trapped for a while inside one cavity formed by the membrane, before an appropriate fluctuation event widens a membrane pore in the right moment so that the object can jump into the next cavity. The process is reflected by a pronounced plateau in the time evolution of the mean squared displacement. Moreover, we investigate the impact of the diffusing object on the deformation of the membrane, which leads to an additional increase in the diffusivity. We think that the described scenario should be directly observable, for instance, in protein diffusion through biological environments.
@@ I Introduction
The diffusive transport of molecules is of interest in biophysics and beyond. Often, this transport happens in porous, heterogeneous environments. Transport processes in biological cells take place in crowded environments [hofling2013anomalous], and diffusion through nanopores has technological applications such as filtration via membranes or in catalysts [wu2020nanoparticle]. Thermal fluctuations do, however, not only lead to the diffusive motion of a molecule, but can also dynamically change the confining environment of this molecule. There have been experimental [sarfati_enhanced_2021], numerical [liNonGaussianNormalDiffusion2019, mohammadi_enhanced_2025], and analytical [marbach_transport_2018, palmieri2012diffusion, reister-gottfried_diffusing_2010] efforts to characterize the effect of a fluctuating environment on a diffusing particle. Depending on the specific setup and the theoretical assumptions, some authors found that the fluctuations of the environment decrease the effective diffusion coefficient [palmieri2012diffusion, reister-gottfried_diffusing_2010], others found that it was increased [sarfati_enhanced_2021, liNonGaussianNormalDiffusion2019, mohammadi_enhanced_2025]. The latter was especially the case when trapping of the diffusing particle in cavities of the porous medium occurred. Fluctuations in a heterogeneus environment can be the cause for non-Gaussian diffusion [liNonGaussianNormalDiffusion2019]. There, the mean-squared displacement (MSD) increases linearly in time, while the displacements do not follow a normal distribution. In many studies, also those mentioned above, the movement is usually limited to a effectively one- or two-dimensional geometry.
@@
Here, we study diffusion in a three-dimensional (3D) fluctuating environment. Examples of porous structures that inherently exist only in three dimensions are cubic membranes. These are complex 3D periodic structures formed by lipid membranes in cell organelles [almsherqi_cubic_2006, almsherqiCubicMembranesMissing2009]. It is also possible to create cubic phases in vitro using amphiphilic molecules.
@@
Mathematically, these structures can be described by periodic surfaces of constant mean curvature (CMC) $H$ [cuiBicontinuousCubicPhases2020]. In the special case of $H=0$, we observe a so-called triply periodic minimal surface (TPMS). CMC surfaces are also obtained by minimizing surface area for a given periodicity while prescribing a fixed ratio of the volumes on the two sides of the surface [grosse-brauckmann_triply_2012]. The TPMS is found in the case of equal volume on both sides of the surface. CMC surfaces can often be thought of as deformed TPMS and are characterized by the TPMS they share their symmetry with. Important examples of TPMSs are the Schwarz P- and D-surfaces, as well as the gyroid (G-surface) discovered by Schoen [von_schnering_how_1987, grosse-brauckmann_triply_2012]. TPMSs and related surfaces separate space into two unconnected volumes that form intertwined channels. The P-surface can be imagined as a cubic lattice of cavities connected by channels in each coordinate direction. That means it forms a network of junctions of six different pipes, see Fig. 1. Similarly the D-surface and the gyroid form channels with junctions of four and three pipes, respectively. Apart from lipid membranes, TPMSs are found in the chitin skeletons of insects, the wings of butterflies, and block-copolymers [hanOverviewMaterialsTriply2018]. They also find technological applications, for example in the design of scaffolds for growing tissue [kapfer2011minimal]. An exact mechanism behind the formation of those structures in biological systems is not known yet. The reason why they are energetically favorable is, however, not that they minimize surface area, as lipid membranes do not have a significant surface tension. Instead, the shape of the membrane molecules dictates a spontaneous curvature $c_{0}$ that is energetically optimal [grosse-brauckmann_triply_2012, schroder2006bicontinuous]. An appropriate energy functional for lipid membranes is therefore the Helfrich free energy [seifertConfigurationsFluidMembranes1997]
@@
where the membrane is described by a surface $S$ and has a bending rigidity $\bar{\kappa}$. A second contribution that depends on an integral of the Gaussian curvature is only relevant if one intends to describe topological changes of the membrane, according to the Gauss-Bonnet theorem [seifertConfigurationsFluidMembranes1997, grosse-brauckmann_triply_2012].
Diffusive motion in the periodic network of channels formed by TPMS has already been investigated in the case of static surfaces [assenzaCurvatureBottlenecksControl2018]. It was found that bottlenecks, if they are present, have an important influence on the transport properties. The dynamics of a diffusing particle then may correspond to a hopping process. It may be trapped in a cavity for a long time before escaping. In fluorescence correlation spectroscopy measurements of the diffusion coefficient of a protein confined in lipid cubic phases, it was found that the impact of the confinement on the diffusion coefficient was less than that of a comparable confinement with hard walls. This suggests that the softness of the confinement and thermal fluctuations play an important role [tanaka2012fluorescence]. We here investigate how diffusion is influenced when a TPMS is undulated by thermal fluctuations, which will lead to a widening and narrowing of the bottlenecks corresponding to the channels formed by the surface. An example for the geometry under investigation is shown in Fig. 1. We focus on the P-surface because its structure is easy to visualize and grasp, which facilitates interpretation of the results. Our approach is, however, straightforward to extend to other CMC surfaces.
@@ II The phase-field method
We use a phase-field approach to track the location of the surfaces. This method describes a surface as the zero-level set of a continuous scalar function, the phase field $\phi(\mathbf{r})$. It is obtained by minimizing a free energy $F\left[\phi\right]$, which is a functional of $\phi$. TPMSs were generated from the phase field model for microemulsions [gozdzTriplyPeriodicSurfaces1996], but this model has many free parameters and could result in a profile of the phase field $\phi$ that is difficult to analyze. Yang et al. [yangPhasefieldApproachMinimizing2010] showed that by minimizing the free energy associated with the Cahn-Hilliard equation starting from appropriate initial conditions, one can generate periodic surfaces of constant mean curvature. We use their approach to generate the initial conditions for our simulations. This free energy is, however, not suited to simulate the dynamic behavior of those surfaces. As explained above, the energy should not depend on the surface area of the membrane. Instead the energy should depend on the mean curvature. Therefore, we use a free energy developed for the modeling of biological membranes and vesicles [duPhaseFieldApproach2004, duPhaseFieldCalculus2011, campeloDynamicModelStationary2006, campeloModelCurvaturedrivenPearling2007, lowengrub_phase-field_2009],
@@
This free energy is designed to describe an interface according to the Helfrich energy in Eq. (1) with $\bar{\kappa}=2\sqrt{2}\epsilon^{3}\kappa/3$ and $\sqrt{2}C_{0}=c_{0}$ [campeloDynamicModelStationary2006]. Consequently, it is minimized by phase fields whose zero-level-set surfaces have constant mean curvature. One obtains a phase field that has a constant value of $\phi=1$ on one side of the surface and $\phi=-1$ on the other side. The parameter $\epsilon$ determines the width of the crossover region between these two domains. For simplicity we focus on minimal surfaces and set $C_{0}=0$ in the remainder of this article.
The following argument by Campelo et al. [campeloDynamicModelStationary2006] on why this energy is related to the mean curvature will be useful in our later analysis. We make the ansatz that $\phi$ is only a function of the distance $d(\mathbf{r})$ to the interface defined by $\phi=0$. By writing $\phi=f\left({d(\mathbf{r})}/{\sqrt{2}\epsilon}\right)$, Eq. (2) becomes
@@
An additional contribution to the free energy arises from the fact that the surface area of the membrane fluctuates around a mean value prescribed by the amount of membrane molecules in the system [campeloDynamicModelStationary2006, duPhaseFieldApproach2004],
@@
This term forces the surface area of the interface to fluctuate around a predefined value set by the parameter $A_{0}$. Here, we set $A_{0}=A[\phi_{0}]$, where $\phi_{0}$ is the initial configuration of the phase field. The value of the parameter $\alpha$ is usually chosen to be very large, so that a constant surface area is enforced. However, in our case, this would hinder all fluctuations of the surface. Instead, we choose an intermediate value of $\alpha$, small enough so that the fluctuations on small length scales are dominated by the bending energy in Eq. (2), but large enough so that the overall periodic structure is not destroyed by thermal fluctuations. Furthermore, the amount of material on both sides of the membrane should be conserved, which leads to a third contribution [duPhaseFieldApproach2004],
@@
We implement the interaction between a diffusing particle and the membrane similar to what had been outlined before [guSimulatingVesicleSubstrate2014]. If an object is moving through the porous structure created by the membrane, there is an energetic cost connected to an overlap of this object and the membrane,
@@
Another option is to use dynamic equations that automatically conserve the integral of $\phi$, eliminating the need for $F_{V}$ [campeloDynamicModelStationary2006, gallenVesicleFormationInduced2023] (Model B in the nomenclature of Halperin, Hohenberg, and Ma [halperinRenormalizationgroupMethodsCritical1974]). They have the disadvantage of producing higher-order spatial derivatives, which makes the numerical approach more involved. Additionally, the stochastic noise would have to be modified to also conserve volume. These conservative dynamics likewise do not capture hydrodynamic interactions. Therefore, we use the simpler dynamics described above. Finally, we note that the (generalized) Fokker-Planck equation for the probability density $P(\phi(\mathbf{r}),\mathbf{R})$, associated with the system of stochastic differential equations, Eqs. (12), has the stationary solution [karmaPhasefieldModelDendritic1999, halperinRenormalizationgroupMethodsCritical1974]
@@ III Membrane fluctuations
For small deviations from the flat surface, the nonlinear terms $F_{A}$ and $F_{V}$ can be neglected. Following Ref. karmaPhasefieldModelDendritic1999, we investigate fluctuations around a flat, rectangular patch of surface $S$. For this purpose, we introduce the coordinates $z$ perpendicular to the surface and $\mathbf{r}_{\perp}$ parallel to the surface. We apply a small perturbation to the surface so that its shape can now be described by the height profile $z=h(\mathbf{r}_{\perp})$. The distance to this perturbed surface is given by
@@
$|S|$ is the surface area. Thus, we recover the known result for the fluctuations of a planar membrane described by the free energy in Eq. (1) (up to a constant factor) [seifertConfigurationsFluidMembranes1997]. Our approach has the advantage that the fluctuation spectrum can be directly predicted from our microscopic equations.
@@ IV Numerical implementation
We measure energies in units of $\kappa_{\mathrm{eff}}=\kappa\epsilon^{3}$, lengths in units of $\epsilon$, and times in units of $M^{-1}_{\phi}\kappa^{-1}$. This means that we effectively set $\epsilon=1$, $\kappa=1$, and $M_{\phi}=1$ in the above equations. Typically, $k_{B}T$ is found in a range from $0.1\kappa_{\mathrm{eff}}$ to $0.01\kappa_{\mathrm{eff}}$ in experiments [monzelMeasuringShapeFluctuations2016]. If not specified otherwise, we set $k_{B}T=0.033$, $\alpha=10$, and $\gamma=0.01$. The dynamical equations are implemented using an Euler scheme, where the linear term in $\delta F_{bend}/\delta\phi$ containing $\nabla^{4}\phi$ is treated implicitly to increase stability. The spatial derivatives are calculated using spectral methods. We set the integration timestep to $\Delta t=0.1$ and $\nu=(\Delta t\Omega)^{-1}$, where $\Omega$ is the volume of the computational domain, to conserve $\int_{\Omega}\phi\ \mathrm{d}^{3}r$ exactly. We discretize $\phi$ on a grid with grid constant $\Delta x=1$.
@@
For the following measurements of diffusion through triply periodic membranes, the phase field describing the initial, unperturbed membrane is generated using the approach of Ref. yangPhasefieldApproachMinimizing2010. The computational domain consists of $3\times 3\times 3$ unit cells of the membrane, where one cell has a length of $d_{c}=40$. A unit cell is therefore about $d_{c}/\sqrt{2}\approx 28$ times larger than the thickness of the transition layer of the phase field from one side of the membrane to the other. This proportion is comparable to cubic membranes found in living systems where structures with unit cells of 50–550 nm can be formed by lipid bilayers with a thickness of less than 10 nm [almsherqiCubicMembranesMissing2009]. For each set of parameters, we simulate 64 independent diffusion processes. Since the presence of the diffusing object influences the membrane, we need to generate a separate realization of the membrane for each realization of the diffusion process. For each realization, we evolve the membrane surface for at least $t_{\mathrm{eq}}=300000$, so that the shape fluctuations equilibrate. This state is saved and used as a starting condition for the subsequent measurements. The starting position of the diffusing object is in the middle of a cavity of the membrane structure.

### v2

@@ Abstract
The transport of individual entities through interconnected structures is a process of practical relevance both in biology and technology. Examples are given by diffusive dynamics of molecules in porous structures. In soft environments, this transport can be strongly influenced by fluctuations of the porous structure itself. Here, we focus on triply periodic membrane structures found both in cell organelles and in synthetic amphiphilic systems. We theoretically study the effect of a complex three-dimensional fluctuating environment on the diffusive motion of a test object, using a phase field approach. The rigid spherical test object is energetically forced to not penetrate the membrane. Generally, the pores of the membrane structure can be smaller than the diffusing object. Yet, fluctuations of the membrane can intermittently widen its pores, still allowing for the motion of the larger particles through them. Thus, the object stays trapped for a while inside one cavity formed by the membrane, before an appropriate fluctuation event widens a membrane pore in the right moment so that the object can jump into the next cavity. The process is reflected by a pronounced plateau in the time evolution of the mean squared displacement. We think that the described scenario should be directly observable, for instance, in protein diffusion through biological environments.
@@ I Introduction
The diffusive transport of molecules is of interest in biophysics and beyond. Often, this transport happens in porous, heterogeneous environments. Transport processes in biological cells take place in crowded environments [1], and diffusion through nanopores has technological applications such as filtration via membranes or in catalysts [2]. Thermal fluctuations do, however, not only lead to the diffusive motion of a molecule, but can also dynamically change the confining environment of this molecule. There have been experimental [3], numerical [4, 5], and analytical [6, 7, 8] efforts to characterize the effect of a fluctuating environment on a diffusing particle. Depending on the specific setup and the theoretical assumptions, some authors found that the fluctuations of the environment decrease the effective diffusion coefficient [7, 8], others found that it was increased [3, 4, 5]. The latter was especially the case when trapping of the diffusing particle in cavities of the porous medium occurred. Fluctuations in a heterogeneus environment can be the cause for non-Gaussian diffusion [4]. There, the mean-squared displacement (MSD) increases linearly in time, while the displacements do not follow a normal distribution. In many studies, also those mentioned above, the movement is usually limited to a effectively one- or two-dimensional geometry.
@@
Here, we study diffusion in a three-dimensional (3D) fluctuating environment. Examples of porous structures that inherently exist only in three dimensions are cubic membranes. These are complex 3D periodic structures formed by lipid membranes in cell organelles [9, 10]. It is also possible to create cubic phases in vitro using amphiphilic molecules.
@@
Mathematically, these structures can be described by periodic surfaces of constant mean curvature (CMC) $H$ [11]. In the special case of $H=0$, we observe a so-called triply periodic minimal surface (TPMS). CMC surfaces are also obtained by minimizing surface area for a given periodicity while prescribing a fixed ratio of the volumes on the two sides of the surface [12]. The TPMS is found in the case of equal volume on both sides of the surface. CMC surfaces can often be thought of as deformed TPMS and are characterized by the TPMS they share their symmetry with. Important examples of TPMSs are the Schwarz P- and D-surfaces, as well as the gyroid (G-surface) discovered by Schoen [13, 12]. TPMSs and related surfaces separate space into two unconnected volumes that form intertwined channels. The P-surface can be imagined as a cubic lattice of cavities connected by channels in each coordinate direction. That means it forms a network of junctions of six different pipes, see Fig. 1. Similarly the D-surface and the gyroid form channels with junctions of four and three pipes, respectively. Apart from lipid membranes, TPMSs are found in the chitin skeletons of insects, the wings of butterflies, and block-copolymers [14]. They also find technological applications, for example in the design of scaffolds for growing tissue [15]. An exact mechanism behind the formation of those structures in biological systems is not known yet. The reason why they are energetically favorable is, however, not that they minimize surface area, as lipid membranes do not have a significant surface tension. Instead, the shape of the membrane molecules dictates a spontaneous curvature $c_{0}$ that is energetically optimal [12, 16]. An appropriate energy functional for lipid membranes is therefore the Helfrich free energy [17]
@@
where the membrane is described by a surface $S$ and has a bending rigidity $\bar{\kappa}$. A second contribution that depends on an integral of the Gaussian curvature is only relevant if one intends to describe topological changes of the membrane, according to the Gauss-Bonnet theorem [17, 12].
Diffusive motion in the periodic network of channels formed by TPMS has already been investigated in the case of static surfaces [18]. It was found that bottlenecks, if they are present, have an important influence on the transport properties. The dynamics of a diffusing particle then may correspond to a hopping process. It may be trapped in a cavity for a long time before escaping. In fluorescence correlation spectroscopy measurements of the diffusion coefficient of a protein confined in lipid cubic phases, it was found that the impact of the confinement on the diffusion coefficient was less than that of a comparable confinement with hard walls. This suggests that the softness of the confinement and thermal fluctuations play an important role [19]. We here investigate how diffusion is influenced when a TPMS is undulated by thermal fluctuations, which will lead to a widening and narrowing of the bottlenecks corresponding to the channels formed by the surface. An example for the geometry under investigation is shown in Fig. 1. We focus on the P-surface because its structure is easy to visualize and grasp, which facilitates interpretation of the results. Our approach is, however, straightforward to extend to other CMC surfaces.
@@ II The phase-field method
We use a phase-field approach to track the location of the surfaces. This method describes a surface as the zero-level set of a continuous scalar function, the phase field $\phi(\mathbf{r})$. It is obtained by minimizing a free energy $F\left[\phi\right]$, which is a functional of $\phi$. TPMSs were generated from the phase field model for microemulsions [20], but this model has many free parameters and could result in a profile of the phase field $\phi$ that is difficult to analyze. Yang et al. [21] showed that by minimizing the free energy associated with the Cahn-Hilliard equation starting from appropriate initial conditions, one can generate periodic surfaces of constant mean curvature. We use their approach to generate the initial conditions for our simulations. This free energy is, however, not suited to simulate the dynamic behavior of those surfaces. As explained above, the energy should not depend on the surface area of the membrane. Instead the energy should depend on the mean curvature. Therefore, we use a free energy developed for the modeling of biological membranes and vesicles [22, 23, 24, 25, 26],
@@
This free energy is designed to describe an interface according to the Helfrich energy in Eq. (1) with $\bar{\kappa}=2\sqrt{2}\epsilon^{3}\kappa/3$ and $\sqrt{2}C_{0}=c_{0}$ [24]. Consequently, it is minimized by phase fields whose zero-level-set surfaces have constant mean curvature. One obtains a phase field that has a constant value of $\phi=1$ on one side of the surface and $\phi=-1$ on the other side. The parameter $\epsilon$ determines the width of the crossover region between these two domains. For simplicity we focus on minimal surfaces and set $C_{0}=0$ in the remainder of this article.
The following argument by Campelo et al. [24] on why this energy is related to the mean curvature will be useful in our later analysis. We make the ansatz that $\phi$ is only a function of the distance $d(\mathbf{r})$ to the interface defined by $\phi=0$. By writing $\phi=f\left({d(\mathbf{r})}/{\sqrt{2}\epsilon}\right)$, Eq. (2) becomes
@@
An additional contribution to the free energy arises from the fact that the surface area of the membrane fluctuates around a mean value prescribed by the amount of membrane molecules in the system [24, 22],
@@
This term forces the surface area of the interface to fluctuate around a predefined value set by the parameter $A_{0}$. Here, we set $A_{0}=A[\phi_{0}]$, where $\phi_{0}$ is the initial configuration of the phase field. The value of the parameter $\alpha$ is usually chosen to be very large, so that a constant surface area is enforced. However, in our case, this would hinder all fluctuations of the surface. Instead, we choose an intermediate value of $\alpha$, small enough so that the fluctuations on small length scales are dominated by the bending energy in Eq. (2), but large enough so that the overall periodic structure is not destroyed by thermal fluctuations. Furthermore, the amount of material on both sides of the membrane should be conserved, which leads to a third contribution [22],
@@
We implement the interaction between a diffusing particle and the membrane similar to what had been outlined before [27]. If an object is moving through the porous structure created by the membrane, there is an energetic cost connected to an overlap of this object and the membrane,
@@
Another option is to use dynamic equations that automatically conserve the integral of $\phi$, eliminating the need for $F_{V}$ [24, 28] (Model B in the nomenclature of Halperin, Hohenberg, and Ma [29]). They have the disadvantage of producing higher-order spatial derivatives, which makes the numerical approach more involved. Additionally, the stochastic noise would have to be modified to also conserve volume. These conservative dynamics likewise do not capture hydrodynamic interactions. Therefore, we use the simpler dynamics described above. Finally, we note that the (generalized) Fokker-Planck equation for the probability density $P(\phi(\mathbf{r}),\mathbf{R})$, associated with the system of stochastic differential equations, Eqs. (12), has the stationary solution [30, 29]
@@ III Membrane fluctuations
For small deviations from the flat surface, the nonlinear terms $F_{A}$ and $F_{V}$ can be neglected. Following Ref. 30, we investigate fluctuations around a flat, rectangular patch of surface $S$. For this purpose, we introduce the coordinates $z$ perpendicular to the surface and $\mathbf{r}_{\perp}$ parallel to the surface. We apply a small perturbation to the surface so that its shape can now be described by the height profile $z=h(\mathbf{r}_{\perp})$. The distance to this perturbed surface is given by
@@
$|S|$ is the surface area. Thus, we recover the known result for the fluctuations of a planar membrane described by the free energy in Eq. (1) (up to a constant factor) [17]. Our approach has the advantage that the fluctuation spectrum can be directly predicted from our microscopic equations.
@@ IV Numerical implementation
We measure energies in units of $\kappa_{\mathrm{eff}}=\kappa\epsilon^{3}$, lengths in units of $\epsilon$, and times in units of $M^{-1}_{\phi}\kappa^{-1}$. This means that we effectively set $\epsilon=1$, $\kappa=1$, and $M_{\phi}=1$ in the above equations. Typically, $k_{B}T$ is found in a range from $0.1\kappa_{\mathrm{eff}}$ to $0.01\kappa_{\mathrm{eff}}$ in experiments [31]. If not specified otherwise, we set $k_{B}T=0.033$, $\alpha=10$, and $\gamma=0.01$. The dynamical equations are implemented using an Euler scheme, where the linear term in $\delta F_{bend}/\delta\phi$ containing $\nabla^{4}\phi$ is treated implicitly to increase stability. The spatial derivatives are calculated using spectral methods. We set the integration timestep to $\Delta t=0.1$ and $\nu=(\Delta t\Omega)^{-1}$, where $\Omega$ is the volume of the computational domain, to conserve $\int_{\Omega}\phi\ \mathrm{d}^{3}r$ exactly. We discretize $\phi$ on a grid with grid constant $\Delta x=1$.
@@
For the following measurements of diffusion through triply periodic membranes, the phase field describing the initial, unperturbed membrane is generated using the approach of Ref. 21. The computational domain consists of $3\times 3\times 3$ unit cells of the membrane, where one cell has a length of $d_{c}=40$. A unit cell is therefore about $d_{c}/\sqrt{2}\approx 28$ times larger than the thickness of the transition layer of the phase field from one side of the membrane to the other. This proportion is comparable to cubic membranes found in living systems where structures with unit cells of 50–550 nm can be formed by lipid bilayers with a thickness of less than 10 nm [10]. For each set of parameters, we simulate 192 independent diffusion processes. Since the presence of the diffusing object influences the membrane, we need to generate a separate realization of the membrane for each realization of the diffusion process. For each realization, we evolve the membrane surface for at least $t_{\mathrm{eq}}=300000$, so that the shape fluctuations equilibrate. This state is saved and used as a starting condition for the subsequent measurements. We generate 64 such starting configurations. The starting position of the diffusing object is in the center of a cavity of the membrane structure. For each starting configuration of the membrane, we simulate three different starting positions of the particle.